OpenStudy (anonymous):

The area of a triangle is 45.5 cm2 and the base is 10 cm. What is the height of the triangle?

OpenStudy (anonymous):

Area of a triangle is \[\large Area~=~ \frac{1}{2} \times Base \times Height\] Just fill in the values of \[\large Area~=~45.5\\\large Base~=~10\] and then solve for the height be getting it alone :)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

What don't you get?

OpenStudy (anonymous):

how to get the base

OpenStudy (anonymous):

i mean width

OpenStudy (anonymous):

We already know that the base is 10. We'r looking for height.

OpenStudy (anonymous):

Here, I'll show you what it looks like. \[\large 45.5~=~\frac{1}{2} \times 10 \times Height\] Get the "Height" value alone.

OpenStudy (anonymous):

yes whats the height i dont know how to find it

OpenStudy (anonymous):

its 9.1